Circular No. 6595

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 46P/WIRTANEN T. Farnham and D. Schleicher, Lowell Observatory, report on narrow-

band photometry of comet 46P with the 0.79-m telescope on Mar. 5.1 UT, at r = 1.07: “Mean Haser-model production rates were log Q(OH) = 27.89, log Q(CN) = 25.43, and log Q(C_{2}) = 25.42, and the mean dust-production rate was log (Afρ) = 2.08; combining these results with those previously obtained with the Hall 1.07-m telescope on Feb. 12.1 and 15.1 (r = 1.14– 1.12 AU) implies that the water production is essentially unchanged over this three-week interval, while the CN and C_{2} production have increased by 50–60 percent and dust production has increased by over a factor of 2. Our measured OH production in mid-Feb. is in good agreement with the water production estimated by Bertaux based on Lyα emission (IAUC 6565).”

K. Jockers, Max Planck Institute for Aeronomy, Katlenburg-Lindau, reports that an H_{2}O^{+} tail was detected in comet 46P with the 2-m telescope of the Center AMEI at Peak Terskol in the Northern Caucasus on Mar. 11.

GS 1843+00 T. Takeshima, Goddard Space Flight Center and Universities Space

Research Association, reports: “The countrate history of the transient x-ray binary pulsar GS 1843+009 obtained with ASM onboard RXTE indicates that the source intensity rose between 1997 Jan. 26 and Feb. 1 to a flux level of 15 mCrab in the band 2–10 keV and remained at 15–30 mCrab level since then. This is the first time that the source has been seen in the x-ray band since the discovery with Ginga in Apr. 1988 (IAUC 4583, 4587). An RXTE observation with PCA and HEXTE was carried out on Mar. 5.0875–5.2549 UT. The averaged source countrate in PCA (2–60 keV) was 806 counts/s, corresponding to a flux level of 62 mCrab. The pulse period (barycentric) was 29.565 ± 0.002 s at the epoch of Mar. 5.1712, which is consistent with the recent BATSE results (IAUC 6586). The obtained pulse period is ∼ 0.2 percent longer than that reported from Ginga observations. The pulse fraction is 31 percent (peak to bottom). The pulse profile is double-peaked, but the secondary minimum is 2.6 percent above the average. The energy spectrum (2–60 keV) is expressed as a sum of an exponential cut-off powerlaw component [photon index −0.20 ± 0.06, cut-off energy 13.0 ± 0.4 keV, and 2–10-keV flux of (1.09 ± 0.12) × 10^{−6} erg s^{−1} cm^{−2}] and a blackbody component [kT = 1.16 ± 0.07 keV and 2–10-keV flux (2.4 ± 0.2) × 10^{−6} erg s^{−1} cm^{−2}] with photoelectric absorption by cold matter of n_{H} = (6.2 ± 0.5) × 10^{22} cm^{−2}.”

1997 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green